# Lightweight Dataset Pruning without Full Training via Example Difficulty and Prediction Uncertainty

### v2

@@ 4.1 Experimental Settings
We assessed the performance of our proposed method in three key scenarios: image classification, image classification with noisy labels and corrupted images. In addition, we validate cross-architecture generalization on three-layer CNN, VGG-16 (Simonyan and Zisserman, 2015), ResNet-18 and ResNet-50 (He et al., 2015).
Hyperparameters. For training CIFAR-10 and CIFAR-100, we train ResNet-18 for 200 epochs with a batch size of 128. SGD optimizer with momentum of 0.9 and weight decay of 0.0005 is used. The learning rate is initialized as 0.1 and decays with the cosine annealing scheduler. As Zhang et al. (2024) show that smaller batch size boosts performance at high pruning rates, we also halved the batch size for 80% pruning, and for 90% we reduced it to one-fourth. For ImageNet-1k, ResNet-34 is trained for 90 epochs with a batch size of 256 across all pruning ratios. An SGD optimizer with a momentum of 0.9, a weight decay of 0.0001, and an initial learning rate of 0.1 is used, combined with a cosine annealing scheduler.
Baselines. The baselines considered in this study are listed as follows [^2] Infomax (Tan et al., 2025) was excluded as it employs different base hyperparameters in the original paper compared to other baselines and does not provide publicly available code. See Appendix A.1 for more discussion.: (1) Random; (2) Entropy (Coleman et al., 2020); (3) Forgetting (Toneva et al., 2018); (4) EL2N (Paul et al., 2021); (5) AUM (Pleiss et al., 2020); (6) Moderate (Xia et al., 2022); (7) Dyn-Unc (He et al., 2024); (8) TDDS (Zhang et al., 2024); (9) CCS (Zheng et al., 2022); and (10) $\mathbb{D}^{2}$ (Maharana et al., 2023). To ensure a fair comparison, all methods were trained with the same base hyperparameters for training, and the best hyperparameters reported in their respective original works for scoring. Technical details are provided in the Section A.1.
Figure 4: Comparison in total time spent on CIFAR datasets.
@@ 4.2 Image Classification Benchmarks
Table 1 presents the test accuracy for image classification results on CIFAR-10 and CIFAR-100. Our pruning method consistently outperforms other baselines, particularly when combined with Beta sampling. While the DUAL score exhibits competitive performance in lower pruning ratios, its accuracy degrades with more aggressive pruning. Our Beta sampling effectively mitigates this performance drop.
Notably, the DUAL score only requires training a single model for only 30 epochs, significantly reducing the computational cost. In contrast, the second-best methods, Dyn-Unc and CCS, rely on scores computed over a full 200-epoch training cycle, making them considerably less efficient. Even considering subset selection, score computation, and subset training, the total time remains less than a single full training run, as shown in Figure 4. Specifically, on CIFAR-10, our method achieves lossless pruning up to a 50% pruning ratio while saving 35.5% of total training time.
Table 2: Comparison of test accuracy on ImageNet-1k. The model trained with the full dataset achieves 73.1% test accuracy. The best result in each pruning ratio is highlighted in bold.
Pruning Rate 30% 50% 70% 80% 90%
@@
We also evaluate our pruning method on the large-scale dataset, ImageNet-1k. The DUAL score is computed during training, specifically at epoch 60, which is 33% earlier than the original train epoch used to compute scores for other baseline methods. As shown in Table 2, Dyn-Unc performs worse than random pruning across all pruning ratios, and we attribute this undesirable performance to its limited total training epochs (only 90), which is insufficient for Dyn-Unc to fully capture the training dynamics of each sample. In contrast, our DUAL score, combined with Beta sampling, outperforms all competitors while requiring the least computational cost. The DUAL score’s ability to consider both training dynamics and the difficulty of examples enables it to effectively identify uncertain samples early in the training process, even when training dynamics are limited. Remarkably, for 90% pruned Imagenet-1K, it maintains a test accuracy of 60.0%, surpassing the previous state-of-the-art (SOTA) by a large margin.
@@ 4.3.1 Label Noise and Image Corruption
To verify this, we evaluate our method by introducing a specific proportion of symmetric label noise (Patrini et al., 2017; Xia et al., 2020; Li et al., 2022) and applying five different types of image corruptions (Wang et al., 2018; Hendrycks and Dietterich, 2019; Xia et al., 2021). We use CIFAR-100 with ResNet-18 and Tiny-ImageNet with ResNet-34 for these experiments. On CIFAR-100, we test label noise and image corruption ratios of 20%, 30%, and 40%. For Tiny-ImageNet, we use a 20% ratio of label noise and image corruption. We prune the label noise-added dataset using a model trained for 50 epochs and the image-corrupted dataset with a model trained for 30 epochs using DUAL pruning—both significantly lower than the 200 epochs used by other methods. For detailed experimental settings, please refer to Appendix A.2.
@@ 4.3.2 Cross-Architecture Generalization
We also evaluate the ability to transfer scores across various model architectures. To be specific, if we can get high-quality example scores for pruning by using a simpler architecture than one for the training, our DUAL pruning would become even more efficient in time and computational cost. Therefore, we focus on the cross-architecture generalization from relatively small networks to larger ones with three-layer CNN, VGG-16, ResNet-18, and ResNet-50. Competitors are selected from each categorized group of the pruning approach: EL2N from difficulty-based, Dyn-Unc from uncertainty-based, and CCS from the geometry-based group.
For instance, we get training dynamics from the ResNet-18 and then calculate the example scores. Then, we prune samples using scores calculated from ResNet-18, and train selected subsets on ResNet-50. The result with ResNet-18 and ResNet-50 is described in Table 3. Surprisingly, the coreset shows competitive performance to the baseline, where the baseline refers to the test accuracy after training a coreset constructed based on the score calculated from ResNet-50. For all pruning cases, we observe that our methods reveal the highest performances. Specifically, when we prune 70% and 90% of the original dataset, we find that all other methods fail, showing worse test accuracies than random pruning.
Table 3: Cross-architecture generalization performance on CIFAR-100 from ResNet-18 to ResNet-50. We report an average of five runs. ‘R50 $\rightarrow$ R50’ stands for score computation on ResNet-50, as a baseline. Test accuracy with full dataset is 80.1%.
@@
Random 74.47 ±$\pm 0.67$ 70.09 ±$\pm 0.42$ 60.06 ±$\pm 0.99$ 41.91 ±$\pm 4.32$
EL2N 76.42 ±$\pm 1.00$ 69.14 ±$\pm 1.00$ 45.16 ±$\pm 3.21$ 19.63 ±$\pm 1.15$
Dyn-Unc 77.31 ±$\pm 0.34$ 72.12 ±$\pm 0.68$ 59.38 ±$\pm 2.35$ 31.74 ±$\pm 2.31$
CCS 74.78 ±$\pm 0.66$ 69.98 ±$\pm 1.18$ 59.75 ±$\pm 1.41$ 41.54 ±$\pm 3.94$
DUAL 78.03 ±$\pm 0.83$ 72.82 ±$\pm 1.46$ 63.08 ±$\pm 2.45$ 33.65 ±$\pm 2.92$
DUAL +$\beta$ 77.82 ±$\pm 0.65$ 73.98 ±$\pm 0.62$ 66.36 ±$\pm 1.66$ 49.90 ±$\pm 2.56$
[…]
We also test the cross-architecture generalization performance with three-layer CNN, VGG-16, and ResNet-18 in Appendix B.3. Even for a simple model like three-layer CNN, we see our methods show consistent performance, as can be seen in Table 13 in Appendix B.3. This observation gives rise to an opportunity to develop some small proxy networks to get example difficulty with less computational cost. Transfer across models with similar capacities, $e.g.$ from VGG-16 to ResNet-18 and vice versa, also supports the verification of cross-architecture compatibility.
@@ 4.4 Ablation Studies
Hyperparameter Analysis.
Figure 6: Left: Varying T with $J=10$ and $c_{\mathcal{D}}=4$. Right: Varying $c_{\mathcal{D}}$ with $T=30$ and $J=10$.
Here, we investigate the robustness of our hyperparameters, $T$, $J$, and $c_{\mathcal{D}}$. We fix $J$ across all experiments, as it has minimal impact on selection, indicating its robustness (Fig 9, Appendix B). In Figure 6, we assess the robustness of $T$ by varying it from 20 to 200 on CIFAR-100. We find that while $T$ remains highly robust in earlier epochs, increasing $T$ degrades generalization performance. This is expected, as larger $T$ overemphasizes difficult samples due to our difficulty-aware selection. Thus, pruning in earlier epochs (from 30 to 50) proves to be more effective and robust. For the $c_{\mathcal{D}}$, we vary it from 3 to 7 and find robustness, especially in the aggressive pruning regime. All results are averaged across three runs.
Beta Sampling Analysis.
Next, we study the impact of our proposed pruning-ratio-adaptive Beta sampling on existing score metrics. We apply our Beta sampling strategy to prior score-based methods, including Forgetting, EL2N, and Dyn-Unc, using the CIFAR10 and CIFAR100 datasets. By comparing our sampling approach with vanilla threshold pruning, which selects only the highest-scoring samples, we observe that prior score-based methods become remarkably comparable to random pruning after Beta sampling is adjusted (see Table 4).
@@
CIFAR-10 CIFAR-100
Method Thresholding $\beta$-Sampling Thresholding $\beta$-Sampling
Ours 54.95 $\pm$0.42 87.09 (+31.51) $\pm$0.36 34.28 $\pm$1.39 54.54 (+20.26) $\pm$0.09
Even adapted for random pruning, our Beta sampling proves to perform well. Notably, EL2N, which performs poorly on its own, becomes significantly more effective when combined with our sampling method. Similar improvements are also seen with Forgetting and Dyn-Unc scores. This is because our proposed Beta sampling enhances the diversity of selected samples in turn, especially when used with example difficulty-based methods. More results conducted for 80% pruning cases are included in the Section B.4.
Additional Analysis.
In addition to the main results presented in this paper, we also conducted various experiments to further explore the effectiveness of our method. These additional results include an analysis of coreset performance under a time budget ($e.g.$ other score metrics are also computed by using training dynamics up to epoch 30) and Spearman rank correlation was calculated between individual scores and the averaged score across five runs to assess the consistency of scores for each sample. Furthermore, additional results in extreme cases, 30% and 40% of label noise and image corruption can be found in Appendix B.
@@ 5 Conclusion
We introduce the Difficulty and Uncertainty-Aware Lightweight (DUAL) score, a novel scoring metric designed for cost-effective pruning. The DUAL score is the first metric to integrate both difficulty and uncertainty into a single measure, and its effectiveness in identifying the most informative samples early in training is further supported by theoretical analysis. Additionally, we propose pruning-ratio-adaptive sampling to account for sample diversity, particularly when the pruning ratio is extremely high. Our proposed pruning methods, DUAL score and DUAL score combined with Beta sampling demonstrate remarkable performance, particularly in realistic scenarios involving label noise and image corruption, by effectively distinguishing noisy samples.
Data pruning research has been evolving in a direction that contradicts its primary objective of reducing computational and storage costs while improving training efficiency. This is mainly because the computational cost of pruning often exceeds that of full training. By introducing our DUAL method, we take a crucial step toward overcoming this challenge by significantly reducing the computation cost associated with data pruning, making it feasible for practical scenarios. Ultimately, we believe this will help minimize resource waste and enhance training efficiency.
@@ A.1 Details on Baseline Implementation
EL2N (Paul et al., 2021) is defined as the error $L2$ norm between the true labels and predictions of the model. The examples with low scores are pruned out. We calculated error norms at epoch 20 from five independent runs, then the average was used for the EL2N score.
Forgetting (Toneva et al., 2018) is defined as the number of forgetting events, where the model prediction goes wrong after the correct prediction, up until the end of training. Rarely unforgotten samples are pruned out.
AUM (Pleiss et al., 2020) accumulates the margin, which means the gap between the target probabilities and the second largest prediction of a model. They calculate the margin at every epoch and then transform it into an AUM score at the end of the training. Here samples with small margins are considered as mislabeled samples, thus data points with small AUM scores are eliminated.
Entropy (Coleman et al., 2020) is calculated as the entropy of prediction probabilities at the end of training, and then the samples that have high entropy are selected into coreset.
@@
TDDS (Zhang et al., 2024) adapts different hyperparameter for each pruning ratio. As they do not provide full information for implementation, we have no choice but to set parameters for the rest cases arbitrarily. The provided setting for (pruning ratio, computation epoch $T$, the length of sliding window $K$) is (0.3, 70, 10), (0.5, 90, 10), (0.7, 80, 10), (0.8, 30, 10), and (0.9, 10, 5) for CIFAR-100, and for ImageNet-(0.3, 20, 10), (0.5, 20, 10), (0.7, 30, 20). Therefore, we set the parameter for CIFAR-10 as the same as CIFAR-100, and for 80%, 90% pruning on ImageNet-1k, we set them as (30, 20), following the choice for 70% pruning.
CCS (Zheng et al., 2022) for the stratified sampling method, we adapt the AUM score as the original CCS paper does. They assign different hard cutoff rates for each pruning ratio. For CIFAR-10, the cutoff rates are (30%, 0), (50%, 0), (70%, 10%), (80%, 10%), (90%, 30%). For CIFAR-100 and ImageNet-1k, we set them as the same as in the original paper.
D2 (Maharana et al., 2023) for $\mathbb{D}^{2}$ pruning, we set the initial node using forgetting scores for CIFAR-10 and CIFAR-100, we set the number of neighbors $k$, and message passing weight $\gamma$ the same as in the original paper.

### v3

@@ 4.1 Experimental Settings
We assessed the performance of our proposed method in three key scenarios: image classification, image classification with noisy labels and corrupted images. In addition, we validate cross-architecture generalization on three-layer CNN, VGG-16 (Simonyan & Zisserman, 2015), ResNet-18 and ResNet-50 (He et al., 2015).
Hyperparameters
For training CIFAR-10 and CIFAR-100, we train ResNet-18 for 200 epochs with a batch size of 128. SGD optimizer with momentum of 0.9 and weight decay of 0.0005 is used. The learning rate is initialized as 0.1 and decays with the cosine annealing scheduler. As Zhang et al. (2024) show that smaller batch size boosts performance at high pruning rates, we also halved the batch size for 80% pruning, and for 90% we reduced it to one-fourth. For ImageNet-1k, ResNet-34 is trained for 90 epochs with a batch size of 256 across all pruning ratios. An SGD optimizer with a momentum of 0.9, a weight decay of 0.0001, and an initial learning rate of 0.1 is used, combined with a cosine annealing scheduler.
Baselines
The baselines considered in this study are listed as follows: [^1] Infomax (Tan et al., 2025) was excluded as it employs different base hyperparameters in the original paper compared to other baselines and does not provide publicly available code. See Appendix A.1 for more discussion. (1) Random; (2) Entropy (Coleman et al., 2020); (3) Forgetting (Toneva et al., 2018); (4) EL2N (Paul et al., 2021); (5) AUM (Pleiss et al., 2020); (6) Moderate (Xia et al., 2022); (7) Dyn-Unc (He et al., 2024); (8) TDDS (Zhang et al., 2024); (9) CCS (Zheng et al., 2022); and (10) $\mathbb{D}^{2}$ (Maharana et al., 2023). To ensure a fair comparison, all methods were trained using a common set of base hyperparameters (e.g., learning rate, batch size, number of epochs), while any method-specific hyperparameters were set to the optimal values reported for each score metric in their respective original works. Technical details are provided in the Section A.1.
@@ 4.2 Image Classification Benchmarks
Table 1 presents the test accuracy for image classification results on CIFAR-10 and CIFAR-100. Our pruning method consistently outperforms other baselines, particularly when combined with Beta sampling. While the DUAL score exhibits competitive performance in lower pruning ratios, its coreset accuracy degrades with more aggressive pruning. Our Beta sampling effectively mitigates this performance drop here.
Notably, the DUAL score requires training a single model for only 30 epochs for computation, significantly reducing the computational cost. In contrast, the second-best methods, Dyn-Unc and CCS, rely on scores computed over 200 epochs—a full training cycle on the original dataset—which makes them significantly less efficient. Even when accounting for subset selection, score computation, and subset training, the total time remains lower than training the full dataset once, as illustrated in Figure 4. Specifically, on CIFAR-10, our method achieves lossless pruning up to a 50% pruning ratio while saving 35.5% of total training time.
We also evaluate our pruning method on the large-scale dataset, ImageNet-1k. The DUAL score is computed during training, specifically at epoch 60, which is 33% earlier than the original train epoch used to compute scores for other baseline methods. As shown in Table 2, Dyn-Unc performs worse than random pruning across all pruning ratios, and we attribute this undesirable performance to its limited total training epochs (only 90), which is insufficient for Dyn-Unc to fully capture the training dynamics of each sample. In contrast, our DUAL score, combined with Beta sampling, outperforms all competitors while requiring the least computational cost. By considering both training dynamics and the difficulty of examples, DUAL can effectively identify uncertain samples early in the training process, even with limited training dynamics than full training. Remarkably, for 90% pruning on ImageNet-1K, it maintains a test accuracy of 60.0%, surpassing the previous state-of-the-art (SOTA) by a large margin.
Table 2: Comparison of test accuracy of DUAL score with existing coreset selection methods using ResNet34 for ImageNet-1K. The model trained with the full dataset achieves 73.1% test accuracy. The best result in each pruning raio is highlighted in bold.
Pruning Rate ($\rightarrow$) 30% 50% 70% 80% 90%
@@
Figure 4: Comparison in total time spent (full dataset training, score estimation, and subset training) on CIFAR datasets. While other methods remain ineffective as they require more than full training, our method achieves a 15.5% time reduction with only 30% pruning, approaching the efficiency of random pruning.
@@ 4.3 Experiments under More Realistic Scenarios
To verify this, we evaluate our method by introducing a specific proportion of symmetric label noise (Patrini et al., 2017; Xia et al., 2020; Li et al., 2022) and applying five different types of image corruptions (Wang et al., 2018; Hendrycks & Dietterich, 2019; Xia et al., 2021). We use CIFAR-100 with ResNet-18 and Tiny-ImageNet with ResNet-34 for these experiments. On CIFAR-100, we test label noise and image corruption ratios of 20%, 30%, and 40% using a model trained for 30 epochs. For Tiny-ImageNet, we use a 20% ratio of label noise and image corruption. We prune the label-noise-added dataset using a model trained for 50 epochs and the image-corrupted dataset with a model trained for 30 epochs using DUAL pruning—both significantly lower than the 200 epochs used by other methods. For detailed experimental settings, please refer to Appendix A.2.
@@ 4.3.2 Cross-architecture generalization
Next, we evaluate the ability to transfer scores across various model architectures. Especially, if we can get high quality example scores for pruning by using a simpler architecture than one for the training, our DUAL pruning would become even more efficient in time and computational cost. Therefore, we focus on the cross-architecture generalization from relatively small networks to larger ones with three-layer CNN, VGG-16, ResNet-18, and ResNet-50. Results are summarized in Table 3.
Competitors are selected from each categorized group of the pruning approach: EL2N from difficulty-based, Dyn-Unc from uncertainty-based, and CCS from the geometry-based group. Standard deviations are omitted here due to space limit; please refer to Appendix B.3 for details.
Table 3: Cross-architecture generalization performance on CIFAR-100 from ResNet-18 to ResNet-50. ‘(R50)’ marker denotes the score is computed on ResNet-50, serving as a baseline. DUAL + $\beta$ means our Beta sampling with DUAL scores. We report an average of five runs. Test accuracy on full dataset with ResNet-50 is 80.1%.
@@
Random 77.17 73.74 66.66 40.48
EL2N 79.46 74.85 58.75 16.19
Dyn-Unc 79.90 75.78 61.75 25.08
CCS 77.24 73.81 66.66 40.31
DUAL 79.48 76.47 68.56 29.82
DUAL + $\beta$ 79.53 75.08 67.54 50.34
DUAL (R50) 79.60 76.64 68.60 29.84
DUAL (R50) + $\beta$ 79.63 76.49 70.37 50.27
Specifically, when pruning 90% of the original dataset, we find that other methods all fail, showing worse test accuracies than random pruning. However, our proposed pruning method consistently shows the powerful ability to generalize across the various network architectures, outperforming or being on par with other computationally expensive baselines. More results are provided in Appendix B.3.
@@ 4.4 Ablation Studies
Hyperparameter Analysis
In this section, we investigate the robustness of our hyperparameters, $T$, $J$, and $c_{\mathcal{D}}$. We fix $J$ across all experiments, as it has minimal impact on selection, indicating its robustness (Figure 9, Appendix B). In Figure 6, we assess the robustness of $T$ by varying it from 20 to 200 on CIFAR-100. We find that while $T$ is highly robust in early epochs, increasing it eventually degrades generalization. This is expected, as larger $T$ overemphasizes difficult samples due to our difficulty-aware selection. Thus, pruning in earlier epochs (30 to 50) proves more effective and robust. For $c_{\mathcal{D}}$, we vary it from 3 to 6 and observe consistent performance, indicating robustness to its choice as well.
Figure 6: Left: T varying while $J=10$ and $c_{\mathcal{D}}=4$. Right: $c_{\mathcal{D}}$ varying while $T=30$ and $J=10$. Three runs are averaged.
Beta sampling with existing scores
Next, we study the impact of our proposed pruning-ratio-adaptive Beta sampling on existing score metrics. We apply our Beta sampling strategy to other score-based methods, including Forgetting, EL2N, and Dyn-Unc, on the CIFAR10 and CIFAR100 datasets. Compared to vanilla threshold pruning, which selects only the highest-scoring samples, we observe that previous methods become competitive when Beta sampling is adjusted (see Table 4). For the case of random pruning combined with Beta sampling, we do not use any score but select samples only with Beta sampling.
@@
CIFAR-10
Method Thresholding $\beta$-Sampling
Random 83.74 $\pm$ 0.21 83.31 (-0.43) $\pm$ 0.14
EL2N 38.74 $\pm$ 0.75 87.00 (+48.26) $\pm$ 0.45
Forgetting 46.64 $\pm$ 1.90 85.67 (+39.03) $\pm$0.13
Dyn-Unc 59.67 $\pm$ 1.79 85.33 (+25.66) $\pm$ 0.20
Ours 54.95 $\pm$ 0.42 87.09 (+32.14) $\pm$ 0.36
CIFAR-100
Method Thresholding $\beta$-Sampling
Random 45.09 $\pm$ 1.26 51.76 (+6.67) $\pm$ 0.25
EL2N 8.89 $\pm$ 0.28 53.97 (+45.08) $\pm$ 0.63
Forgetting 26.87 $\pm$ 0.73 52.40 (+25.53) $\pm$ 0.43
Dyn-Unc 34.57 $\pm$ 0.69 51.85 (+17.28) $\pm$ 0.35
Ours 34.28 $\pm$ 1.39 54.54 (+20.26) $\pm$ 0.09
Even with random pruning, our Beta sampling continues to perform well. Notably, EL2N, which performs poorly on its own, becomes significantly more effective when combined with our sampling method. Similar improvements are also seen with Forgetting and Dyn-Unc scores. This is because our proposed Beta sampling enhances the diversity of selected samples, especially when used with example difficulty-based methods. More results conducted at 80% are included in the Section B.6.
Additional Analysis
In addition to the main results presented in this paper, we conducted various experiments to validate the effectiveness of our method, which is provided in Appendix B. For instance, these investigations include: (i) calculating the Spearman rank correlation between individual DUAL scores and their average DUAL score across five runs to assess score consistency; and (ii) analyzing coreset performance under a time budget. These analyses are presented in Figure 7 and Figure 8, respectively, in Appendix B.
Furthermore, detailed results on extreme cases of label noise (ranging from 20% to 40% for CIFAR-100 and 20% for Tiny-ImageNet) are presented in Section B.1. Similar comprehensive results for various image corruptions can be found in Section B.2. The generalization performance of our method across other network architectures is further detailed in Section B.3. Additionally, results for long-tailed data classification using the CIFAR-10-LT and CIFAR-100-LT datasets are provided in Section B.4. Lastly, a comparison with dynamic pruning methods such as Qin et al. (2024) and Yuan et al. (2025) is provided in Section B.5.
@@ 5 Conclusion
This paper introduces Difficulty and Uncertainty-Aware Lightweight (DUAL), a novel scoring metric for cost-effective pruning. DUAL is the first metric to combine difficulty and uncertainty into a single measure, and its effectiveness in identifying the most informative samples early in training is further supported by theoretical analysis. Further, we propose pruning-ratio-adaptive sampling to consider the sample diversity when the pruning ratio is extremely high. Our proposed DUAL score, combined with Beta sampling, shows remarkable performance, particularly under label noise and image corruption by effectively distinguishing noisy samples. Future work could explore extending this approach to unsupervised settings.
Acknowledgement
This work was supported by two Institute of Information & communications Technology Planning & Evaluation (IITP) grants (No. RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST); No. RS-2024-00457882, National AI Research Lab Project) funded by the Korean government (MSIT), and a National Research Foundation of Korea (NRF) grant (No. NRF-2019R1A5A1028324) funded by the Korean government (MSIT).
Impact Statement
This paper presents work on data pruning to advance machine learning, with the potential for positive societal impact through improved efficiency.
@@ A.1 Details on Baseline Implementation
EL2N (Paul et al., 2021) is defined the error L2 norm between the true labels and predictions of model. Then examples with low scores are pruned out. We calculate error norm at epoch 20 from five independent runs, then the average was used for EL2N score.
Forgetting (Toneva et al., 2018) is defined as the number of forgetting events, where the model prediction goes wrong after the correct prediction, up until the end of training. Rarely are unforgotten samples pruned out.
AUM (Pleiss et al., 2020) accumulates the margin, which means the gap between the target probabilities and the second largest prediction of model. They calculate the margin at every epoch and then transform it into an AUM score at the end of the training. Here samples with small margin are considered as mislabeled samples, thus data points with small AUM scores are eliminated.
Entropy (Coleman et al., 2020) is calculated as the entropy of prediction probabilities at the end of training, then the samples which have high entropy are selected into coreset.
@@
TDDS (Zhang et al., 2024) adapts different hyperparameter for each pruning ratio. As they do not provide full information for implementation, we have no choice but set parameters for the rest case arbitrarily. The provided setting for (pruning ratio, computation epoch $T$, the length of sliding window $K$) is (0.3, 70, 10), (0.5, 90, 10), (0.7, 80, 10), (0.8, 30, 10), and (0.9, 10, 5) for CIFAR-100, and for ImageNet-(0.3, 20, 10), (0.5, 20, 10), (0.7, 30, 20). Therefore, we set the parameter for CIFAR-10 as the same with CIFAR-100, and 80%, 90% pruning on ImageNet-1K, we set them as (30, 20), following the choice for 70% pruning.
CCS (Zheng et al., 2022) for stratified sampling method, we adapt AUM score as the original CCS paper does. They assign different hard cutoff rate for each pruning ratio, For CIFAR10, the cutoff rate is (30%, 0), (50%, 0), (70%, 10%), (80%, 10%), (90%, 30%). For CIFAR100 and ImageNet-1K, we set them as the same with the original paper. As explicitly mentioned in Appendix B of Zheng et al. (2022), we use the AUM score calculated at the end of training. This means scores are computed at epoch 200 for the CIFAR-10/100 datasets and at epoch 90 for the ImageNet-1K dataset.
D2 (Maharana et al., 2023) for $\mathbb{D}^{2}$ pruning, we set the initial node using forgetting scores for CIFAR-10 and CIFAR-100, we set the number of neighbors $k$, and message passing weight $\gamma$ as the same with the original paper.
Remark.
As detailed by Zhang et al. (2024) (Section 5.2), TDDS employs 90 epochs for initial full-dataset training. Subsequently, an exhaustive search is conducted to determine an optimal epoch for score computation (e.g., 30 epoch for pruning ImageNet by 70-90%). In our evaluations, we utilized these reported optimal epochs for TDDS across all pruning ratios. While this approach leads to a shorter score computation period for TDDS (excluding the significant overhead of the exhaustive search itself), it is crucial to note that our proposed method consistently achieves significantly higher test accuracy. This advantage is demonstrated by both their reported results and our reproduced experiments.